Suppose you are in charge of a toll bridge that costs essentially nothing to operate. The demand for bridge crossings Q is given by P = 15 – 0.5Q. Initially you charge a toll at $7. Now you reduce the toll to $6. How much consumer surplus can the consumers of the bridge gain from this reduction?
Answer #1

A I5 O Sa /S-P 30 - 2P 7 30-207 6 - 2 C6) 30 l6

Initial CS=AED

AED=(15-7)*16/2

AED=64

CS after fall in price.

New CS=ABC

ABC=(15-6)*18/2

ABC=81

Gain in CS=81-64=17